OpenStudy (anonymous):

Which of the following is the conjugate of a complex number with 6 as the real part and -5 as the imaginary part? 6 - 5i -6 - 5i 6 + 5i -6 + 5i

OpenStudy (anonymous):

@SiennaSurfer @surjithayer @ParthKohli

OpenStudy (solomonzelman):

a and b are constants (or negative/positive intergers) and for any "a" and "b" , a-bi conjugate = a+bi a+bi conjugate = a-bi -a+bi conjugate = -a-bi -a-bi conjugate = -a+bi

OpenStudy (anonymous):

so the answer would be 6+5i @SolomonZelman

OpenStudy (solomonzelman):

Yeah!!!!

OpenStudy (solomonzelman):

Do you know what the purpose of this is?

OpenStudy (anonymous):

thank you so much, i understand it now.

OpenStudy (solomonzelman):

Do you know what the purpose of this is?

OpenStudy (solomonzelman):

the purpose of conjugate I mean

OpenStudy (anonymous):